OpenStudy (anonymous):

Write a polynomial function in standard form with the given zeros. x= 1,1,2 Steps please?

OpenStudy (anonymous):

first you have to set all three factors up to equal 0

OpenStudy (anonymous):

x = 1 x-1 = 0

OpenStudy (anonymous):

since the zeros have to be at 1 and 2 (with a double root at 1), this must at least be a third degree polynomial... y = (x-1)(x-1)(x-2) multiply out the right side and write it in descending order with the highest degree first down to the least.

OpenStudy (anonymous):

x= 2 x-2 = 0

OpenStudy (anonymous):